# Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Latent Dynamics Model

### v1

@@ Abstract
The manipulation of deformable objects by robotic systems presents a significant challenge due to their complex and infinite-dimensional configuration spaces. This paper introduces a novel approach to Deformable Object Manipulation (DOM) by emphasizing the identification and manipulation of Structures of Interest (SOIs) in deformable fabric bags. We propose a bimanual manipulation framework that leverages a Graph Neural Network (GNN)-based latent dynamics model to succinctly represent and predict the behavior of these SOIs. Our approach involves constructing a graph representation from partial point cloud data of the object and learning the latent dynamics model that effectively captures the essential deformations of the fabric bag within a reduced computational space. By integrating this latent dynamics model with Model Predictive Control (MPC), we empower robotic manipulators to perform precise and stable manipulation tasks focused on the SOIs. We have validated our framework through various empirical experiments demonstrating its efficacy in bimanual manipulation of fabric bags. Our contributions not only address the complexities inherent in DOM but also provide new perspectives and methodologies for enhancing robotic interactions with deformable objects by concentrating on their critical structural elements. Experimental videos can be obtained from https://sites.google.com/view/bagbot
@@
Deformable object manipulation, structure of interest, latent dynamics model, bimanual manipulation.
@@ I Introduction
Deformable object manipulation (DOM) [1, 2, 3] is a fundamental capability for robots to meaningfully interact with the physical world and assist in various human tasks. However, the manipulation of deformable objects such as cloth [4], rope [5], and food ingredients[6] is particularly challenging due to their infinite-dimensional configuration space and complex dynamics. Traditional methods in DOM often resort to simplified physics models or data-driven modeling with handcrafted features, which lack adaptability for the varied shapes and dynamics of these objects [7]. Moreover, most current DOM works focus on manipulating the entire object, neglecting the critical structures, i.e., Structures of Interest (SOI), that are essential for subsequent manipulation steps.
Figure 1: (Left) SOI Examples for different deformable object manipulation tasks, e.g., garment hanging, robot-assistive dressing. (Right) Conceptual representation of the manifold with boundary. The manifold encompasses $\operatorname{Int}\mathcal{M}$ and $\partial\mathcal{M}$, where the local neighborhoods of points in $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$ are homeomorphically equivalent to $\operatorname{Int}\mathbb{H}^{n}$ and $\partial\mathbb{H}^{n}$.
@@
In this paper, we introduce the concept of SOI into the realm of DOM (see Fig. 1 for an example), a paradigm shift that emphasizes the importance of identifying and manipulating key structural components rather than the entire object. This focus on the SOI is motivated by the observation that successful DOM tasks typically involve the manipulation of these key areas. By targeting these SOIs, we can reduce the computational load significantly, as modeling the complete 3D dynamics of the deformable object is unnecessary and burdensome for the task at hand.
Furthermore, this work is pioneering in considering the bimanual manipulation of a deformable fabric bag as shown in Fig. 2. By proposing a novel bimanual manipulation framework using a GNN-based latent dynamics model, we address the complexities of DOM with a focus on SOIs. Our approach is designed to extract the SOI from the observed object point cloud, construct a graph representation, and learn the latent dynamics model that effectively captures the object’s deformations within a compact space. Integrating this model with model predictive control (MPC), we enable robots to achieve accurate and stable manipulation of deformable bags.
The main contributions of this work include:
  • Pioneering the bimanual manipulation of deformable fabric bags by focusing on SOIs.
  • Introducing the SOI concept for representing deformable object states, which emphasizes the manipulation of critical structures.
  • Designing a GNN-based method to learn a latent dynamics model from partial point cloud data, particularly focusing on the SOIs.
  • Implementing MPC based on the latent dynamics for generating optimal manipulation actions centered around SOIs.
Various empirical experiments on the bimanual manipulation of a fabric bag validate the efficacy of our proposed framework. Our work offers new insights and methodologies for improving robotic systems’ capability for intelligent deformable object manipulation, with a specialized emphasis on the pivotal SOIs.
Figure 3: Bimanual bag manipulation is formulated as a POMDP problem, where the SOI-related observation $\hat{o}_{t}$ is extracted from the original observation $o_{t}$ and governed by $f_{\theta_{\mathrm{dym}}}$.
@@ II Related Work
Deformable object manipulation has been an active research area in robotics. Existing methods can be categorized into model-based and data-driven approaches. Model-based methods rely on simplified physics models to represent deformable objects. Early works used mass-spring models (MSM) to simulate deformation [8, 9]. The finite element method (FEM) provides more accurate modeling of continuum mechanics [10, 11, 12]. However, analytical models require extensive manual tuning and generalization across different materials or shapes remains difficult. Data-driven methods [7, 13, 14] aim to learn models directly from data. Vision-based methods extract geometric features from visual observations to infer deformations [15, 16]. Recent works utilized deep learning on point cloud data and achieved improved modeling accuracy [17]. However, they depend heavily on large labeled datasets. Self-supervised methods were proposed to learn from physical interactions [18, 19]. However, they focused on planar objects and could not handle complex deformations.
Our work is mostly related to robot manipulation using graph neural networks (GNNs). GNNs have shown promising results in learning physics simulations [20, 21, 22, 23, 24] and control policies [25, 26, 27, 28]. Recently, GNNs were introduced to model rope and cloth manipulation [29, 30, 31]. Different from these works, we propose the structure of interest concept to focus on key object components and employ GNNs to learn latent dynamics models for describing complex deformable object behaviors. The integration of the latent dynamics model with MPC also distinguishes our framework from prior arts. In summary, our work aims to advance existing deformable object manipulation by introducing an SOI-based modeling approach. The adoption of state-of-the-art deep learning techniques allows better generalizability across different materials and shapes. The experiments on the bimanual manipulation of fabric bags represent challenging test cases and validate the feasibility of our framework.
@@ III Problem Statement
Given that individual image and depth observations generally do not fully disclose the state of the environment, we approach the task of bimanual bag manipulation as a Partially Observable Markov Decision Process (POMDP) as depicted in Fig. 3. This is formally defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},\mathcal{B},\mathcal{R},\gamma)$, where the state at time $t$, denoted by $s_{t}$, belongs to the state space $\mathcal{S}$ and is not directly observable. The state encapsulates the configuration of the robots and the manipulated object. The corresponding observation at time $t$, denoted by $o_{t}$, is within the observation space $\mathcal{O}$. The state transition model $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$ describes the probability of transitioning from the current state $s_{t}$ to a new state $s_{t+1}$ upon taking an action $a_{t}$ from the action space $\mathcal{A}$, which consists of the combined left and right robotic actions, represented by the Cartesian product $\mathcal{A}=\mathcal{A}^{1}\times\mathcal{A}^{2}$. The function $\mathcal{B}(o_{t}\mid s_{t},a_{t-1})$ specifies the likelihood of observing $o_{t}$ after executing action $a_{t-1}$ and transitioning to state $s_{t}$. The reward function $\mathcal{R}(s_{t},a_{t})$ assigns a valued reward to each state-action pair, and the discount factor $\gamma\in[0,1)$ quantifies the preference for immediate rewards over future rewards.
@@
The goal of this work is to use two 3D-printed robot grippers to grasp the handles to manipulate the opening rim to achieve a target state $\mathbf{g}$. We assume this deformable bag manipulation task is a quasi-static manipulation and dynamic manipulation motions are not considered. At time step $t$, the robots apply action $(\mathbf{a}^{1}_{t},\mathbf{a}^{2}_{t})\in\mathcal{A}$ upon the bag, and we can partially observe transitions of the bag from $\mathbf{o}_{t}$ to $\mathbf{o}_{t+1}$ under the unknown state transitions from $\mathbf{s}_{t}$ to $\mathbf{s}_{t+1}$. However, a complete observation of the bag is not necessary, in our task, the opening rim of the bag is critical for successful manipulation tasks since it not only determines the manipulation task goals but also provides the most informative sensory feedback, such as visual landmarks, during manipulation. We define the Structure of Interest (SOI) in the context of Deformable Object Manipulation (DOM) refers to specific regions or features of a deformable object that are critical for successful manipulation tasks (see Fig. 1 for examples.). Therefore, in this task, we consider the opening rim of the manipulated bag as our SOI points, and topologically, we can define this loop-like structure as a manifold with boundary. As illustrated in Fig. 1, we also define its interior and boundary as $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$, whose points’ neighborhoods are respectively homeomorphic to $\operatorname{Int}\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}>0\}$ and $\partial\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}=0\}$.
@@
With an appropriate perception module, the observation of the Structure of Interest (SOI), denoted by $\hat{o}_{t}$, can be extracted from the overall observation of the bag $o_{t}$. Our approach is based on the insight that it is more efficacious to predict the dynamics of the SOI rather than the entire complex dynamics of the bag. To this end, we employ a Graph Neural Network (GNN) to establish a dynamics model $f_{\theta_{\text{dym}}}$ that is dedicated to learning the transition functions of the SOI, defined as $f_{\theta_{\text{dym}}}:\hat{\mathcal{O}}\times\mathcal{A}\rightarrow\hat{\mathcal{O}}$. This dynamics model accepts as input a sequence of SOI observations $\mathbf{\hat{o}}_{t-n,\ldots,t}\in\mathcal{\hat{O}}$ and actions $(\mathbf{a}^{1}_{t-n,\ldots,t},\mathbf{a}^{2}_{t-n,\ldots,t})\in\mathcal{A}$, and predicts the subsequent observation $\mathbf{\hat{o}}_{t+1}$, where $n$ represents the length of the observation history before the current time step $t$. With the dynamics model, we proceed to cast the bimanual manipulation of the bag as a task within the Model Predictive Control (MPC) framework. Within this MPC setup, the cost function $\mathcal{J}$ quantifies the difference between the final SOI feature points at time step $T$ and the targeted SOI state $\mathbf{g}$. Details on the precise structure of the cost function $\mathcal{J}$ are illustrated in the following section. This cost is minimized to yield an optimal sequence of actions across a temporal horizon of $T$ steps:
$\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:{T-1}}$ $=\underset{\langle\mathbf{a}^{0},\mathbf{a}^{1}\rangle_{0:T-1}\in\mathcal{A}}{\arg\min}\mathcal{J}(\mathbf{\hat{g}},\mathbf{g})$ (1)
$\text{where}\quad\mathbf{\hat{g}}$ $=f_{\theta_{dym}}\left(\mathbf{\hat{o}}_{0},\langle\mathbf{a}^{0},\mathbf{a}^{1}\rangle_{0:{T-1}}\right)$
where $\hat{\mathbf{g}}$ represents the predicted SOI state at time step $T$. Fig. 4 shows the overall framework of our proposed latent dynamics model for the bimanual bag manipulation task.
@@
Figure 4: The conceptual representation of the proposed framework for bimanual deformable fabric bag manipulation is based on the latent SOI dynamics model.
Figure 5: The proposed SOI global particle sampling process. Left: the reconstructed point cloud for the fabric bag. Right: (a) raw SOI point cloud. (b) preprocessed SOI point cloud. (c) Reconstructed SOI Surface. (d) Resampled SOI particles.

### v2

# Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Neural Dynamics Model

@@ Abstract
The manipulation of deformable objects by robotic systems presents significant challenges due to their complex dynamics and infinite-dimensional configuration spaces. This paper introduces a novel approach to deformable object manipulation (DOM) by emphasizing the introduction and manipulation of structures of interest (SOIs) in deformable fabric bags. We propose a bimanual manipulation framework that leverages a graph neural network (GNN)-based neural dynamics model to succinctly represent and predict the behavior of these SOIs. Our approach involves global particle sampling process to construct a particle representation from partial point clouds of the SOIs and learning the neural dynamics model that effectively captures the essential deformations of the SOIs for fabric bags. By integrating this neural dynamics model with model predictive control (MPC), we enable robotic manipulators to perform precise and stable manipulation tasks focused on the SOIs. We validate our new framework through various experiments that demonstrate its efficacy in manipulating deformable bags and T-shirts. Our contributions not only address the complexities inherent in DOM but also provide new perspectives and methodologies for enhancing robotic interactions with deformable materials by concentrating on their critical structural elements. Videos of the conducted experiments can be seen at https://sites.google.com/view/bagbot.
@@
Deformable object manipulation, structure of interest, neural dynamics model, bimanual manipulation.
@@ I Introduction
Deformable object manipulation (DOM) [1, 2, 3] is a fundamental capability for robots to meaningfully interact with the physical world and assist in various human tasks. However, the manipulation of deformable objects such as cloths [4], ropes [5], and food materials [6] is particularly challenging due to their infinite-dimensional configuration space and complex dynamics. Traditional methods in DOM often resort to simplified physics models or data-driven modeling with handcrafted features, which lack adaptability for the varied shapes and dynamics of these objects. Moreover, most current DOM works focus on manipulating the entire object, neglecting the critical structures, i.e., structures of interest (SOIs), that are essential for subsequent manipulation steps.
Figure 1: (Left) SOI Examples for different deformable object manipulation tasks, e.g., garment hanging, and robot-assistive dressing. (Right) Conceptual representation of the manifold with boundary. The manifold encompasses $\operatorname{Int}\mathcal{M}$ and $\partial\mathcal{M}$, where the local neighborhoods of points in $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$ are homeomorphically equivalent to $\operatorname{Int}\mathbb{H}^{n}$ and $\partial\mathbb{H}^{n}$.
@@
In this paper, we introduce the concept of SOI into the realm of DOM (see Fig. 1 for an example), a paradigm shift that emphasizes the importance of leveraging and manipulating key structural components rather than the entire object. This focus on the SOI is motivated by the observation that successful DOM tasks typically involve the manipulation of these key areas. By targeting these SOIs, we can reduce the computational load significantly, as modeling the complete 3D dynamics of the deformable object is unnecessary and burdensome for the task at hand.
In this work, we address the bimanual manipulation of a deformable fabric bag as illustrated in Fig. 2. Our proposed framework introduces a GNN-based neural dynamics model that enables to tackle the challenges associated with deformable object manipulation (DOM), with an emphasis on Structures of Interest (SOIs). The framework seeks to effectively represent the state of the object by extracting SOIs’ particles from the point cloud data, and learning the underlying dynamics within a condensed particle space. By integrating this method with model predictive control (MPC), we enable robots to achieve accurate and stable manipulation of deformable bags. The contributions of this work are as follows:
  • The development of an original bimanual manipulation strategy for deformable fabric bags with an emphasis on the utility of SOIs.
  • The adoption of the SOI concept for deformable object representations using the proposed global particle sampling method, highlighting the significance of key structural elements during manipulation.
  • A new approach to learning the SOI-based neural dynamics models via GNNs, applied to particle sets of SOIs.
  • The application of MPC informed by the neural dynamics model to guide the generation of optimal manipulation actions that prioritize SOIs.
Our experiments demonstrate the potential of this framework in the context of bimanual manipulating deformable objects, such as fabric bags, T-shirts. While our findings suggest improvements in the robotic manipulation of deformable objects by concentrating on SOIs, we acknowledge the ongoing need for research in this area. This work contributes to the broader understanding of intelligent manipulation strategies and opens pathways for future innovations in the robotic handling of complex materials.
@@ II Related Work
Deformable object manipulation (DOM) has been an active research area in robotics. Existing methods can be categorized into model-based and data-driven approaches. Model-based methods rely on simplified physics models to represent deformable objects. Early works used mass-spring models (MSM) to simulate deformation [7]. The finite element method (FEM) provides more accurate modeling of continuum mechanics [8, 9]. However, analytical models require extensive manual tuning and generalization across different materials or shapes remains difficult. Data-driven methods [10, 11] aim to learn models directly from data. Vision-based methods extract geometric features from visual observations to infer deformations [12, 13]. Recent works utilized deep learning on point cloud data and achieved improved modeling accuracy [14]. However, they depend heavily on large labelled datasets. Self-supervised methods were proposed to learn from physical interactions [15, 16]. However, they focused on planar objects and could not handle complex deformations.
Recent advancements in DOM have leveraged graph neural networks (GNNs) to model the complex interactions within varying materials and shapes [17, 18, 19, 20]. Wang et al. [21] proposed an offline-online learning framework that utilizes GNNs for the deformation model in cable manipulation. This method captures the intricate physics of cables, achieving a balance between offline learning from a rich dataset and online refinement through interaction. Deng et al. [22] introduced a deep reinforcement learning approach that employs local GNNs for a goal-conditioned rearranging of deformable objects. Their method efficiently adapts to the changing dynamics of the objects during manipulation tasks. In the domain of elasto-plastic object manipulation, Shi et al. [23] presented RoboCraft, which combines perception and simulation through GNNs to understand and shape 3D objects. Furthermore, the GDOOM framework by Ma et al. [24] focuses on learning latent graph dynamics for visual manipulation, offering a comprehensive approach to predicting deformable object behavior.
In the specific area of deformable bag manipulation, a number of approaches have been proposed. Xu et al. [10] developed DextAIRity, a system that simplifies the manipulation of deformable objects by harnessing pneumatic controls. Gu et al. [25] introduced ShakingBot, which demonstrates dynamic manipulation skills for bagging tasks, highlighting the potential to handle high-speed and complex deformable object interactions. Chen et al. [26] focused on the challenge of opening plastic bags and inserting objects, a task with significant practical applications in automation. Lastly, Weng et al. [27] explored the concept of interactive perception for deformable object manipulation, where the perception and interaction are tightly integrated to enhance manipulation capabilities.
Our work distinguishes itself from the existing literature by introducing a novel bimanual manipulation framework for deformable fabric bags that places a strong emphasis on structures of interest (SOIs). Unlike previous work that focused on the manipulation of entire objects or specific applications such as fabric or plastic bag handling, our approach centers on identifying and manipulating the SOIs within fabric bags to achieve precise and efficient DOM. We propose a GNN-based latent dynamics model tailored to represent and predict the behavior of these SOIs, which is a subtle yet meaningful advancement from the general-purpose models in the current literature. Furthermore, by integrating this model with model predictive control (MPC), our framework enables accurate and stable manipulation of fabric bags by robotic systems in a computationally efficient manner. Our contributions lie in the specialized focus on SOIs and the combination of GNN-based modeling with MPC, subtly enriching the landscape with fresh viewpoints for enhancing robotic DOMs.
Figure 3: (Left) Conceptual representation of SOI-based bimanual deformable object manipulation problem; (Right) The problem is formulated as a POMDP setting, where the SOI-related observation $\mathbf{o}^{soi}_{t}$ is extracted from the original observation $o_{t}$ and governed by $f_{\theta_{\mathrm{dyn}}}$.
@@ III Problem Statement
Given that individual image and depth observations generally do not fully disclose the state of the environment, we approach the task of bimanual bag manipulation as a Partially Observable Markov Decision Process (POMDP) as depicted in Fig. 3. This is formally defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},\Omega,\mathcal{R},\gamma)$, where the state at time $t$, denoted by $s_{t}$, belongs to the state space $\mathcal{S}$ and is not directly observable. The state encapsulates the configuration of the robots and the manipulated deformable object. The corresponding observation at time $t$, denoted by $o_{t}$, is within the observation space $\mathcal{O}$. The state transition model $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$ describes the probability of transitioning from the current state $s_{t}$ to a new state $s_{t+1}$ upon taking an action $a_{t}$ from the action space $\mathcal{A}$, which consists of the combined left and right robotic actions, represented by the Cartesian product $\mathcal{A}=\mathcal{A}^{0}\times\mathcal{A}^{1}$. The function $\Omega(o_{t}\mid s_{t},a_{t-1})$ specifies the likelihood of observing $o_{t}$ after executing action $a_{t-1}$ and transitioning to state $s_{t}$. The reward function $\mathcal{R}(s_{t},a_{t})$ assigns a valued reward to each state-action pair, and the discount factor $\gamma\in[0,1)$ quantifies the preference for immediate rewards over future rewards.
@@
The goal of this work is to use two robot manipulators to firmly grasp and manipulate the SOI of the deformable object to achieve a target SOI state $\mathbf{y}_{*}$. We assume this deformable object manipulation task is a quasi-static manipulation and dynamic manipulation motions are not considered. As shown in Fig. 3, at time step $t$, the dual-arm manipulators apply action $\{\mathbf{a}^{0}_{t},\mathbf{a}^{1}_{t}\}=\mathbf{a}_{t}\in\mathcal{A}$ upon the deformable object, and we can partially observe transitions of the object from $\mathbf{o}_{t}$ to $\mathbf{o}_{t+1}$ under the unknown state transitions from $\mathbf{s}_{t}$ to $\mathbf{s}_{t+1}$. However, a complete observation of the bag is not necessary, in our task, the opening rim of the bag is critical for successful manipulation tasks since it not only determines the manipulation task goals but also provides the most informative sensory feedback, such as visual landmarks, during manipulation. We define the structure of Interest (SOI) in the context of deformable object manipulation (DOM) refers to specific regions or features of a deformable object that are critical for successful manipulation tasks (e.g., examples in Fig. 1). Therefore, in this task, we consider the opening rim of the manipulated bag as our SOI points, and topologically, we can define this loop-like structure as a manifold with boundary [lee2012smooth]. As illustrated in Fig. 1, we also define its interior and boundary as $\operatorname{Int}\mathcal{M}$ and ${\partial\mathcal{M}}$, whose points’ neighborhoods are respectively homeomorphic to $\operatorname{Int}\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}>0\}$ and $\partial\mathbb{H}^{n}=\{(x_{1},\ldots,x_{n})\mid x_{n}=0\}$.
@@
With an appropriate perception module, the observation of the SOI, denoted by $\mathbf{o}^{soi}_{t}$, can be extracted from the overall observation of the bag $\mathbf{o}_{t}$. Our approach is based on the insight that it is more efficient to predict the dynamics of the SOI rather than the entire complex dynamics of the bag. To this end, we employ a graph neural network (GNN) to establish a dynamics model $f_{\theta_{\text{dyn}}}$ that is dedicated to learning the transition functions of the SOI, defined as $f_{\theta_{\text{dyn}}}:\mathcal{O}^{soi}\times\mathcal{A}\rightarrow\mathcal{O}^{soi}$. This dynamics model accepts as input a sequence of SOI observations $\mathbf{o}^{soi}_{t-n:t}\in\mathcal{O}^{soi}$ and bimanual actions $\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{t-n:t}\subseteq\mathcal{A}$, and predicts the subsequent SOI observation $\mathbf{o}^{soi}_{t+1}$, where $n$ represents the length of the observation history before the current time step $t$. With the dynamics model, we proceed to cast the bimanual manipulation of the bag as a task within the model predictive control (MPC) framework. Within this MPC setup, the cost function $\mathcal{J}$ quantifies the difference between the final SOI observation at time step $T$ and the targeted SOI state $\mathbf{o}^{soi}_{*}$. Details on the precise structure of the cost function $\mathcal{J}$ are illustrated in the following section. This cost is minimized to yield an optimal sequence of actions across a temporal horizon of $T$ steps:
$\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{0:{T-1}}$ $=\underset{\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{0:T-1}\subseteq\mathcal{A}}{\arg\min}\mathcal{J}(\mathbf{o}^{soi}_{T},\mathbf{o}^{soi}_{*})$ (1)
$\text{where}\quad\mathbf{o}^{soi}_{T}$ $=f_{\theta_{dyn}}\left(\mathbf{o}^{soi}_{0},\{\mathbf{a}^{0},\mathbf{a}^{1}\}_{0:{T-1}}\right)$
where $\mathbf{o}^{soi}_{T}$ represents the predicted SOI state at time step $T$.
@@
Figure 4: The conceptual representation of the proposed framework for bimanual deformable fabric bag manipulation using the latent SOI dynamics model.
Figure 5: The proposed global particle sampling process: (a) Reconstructed point cloud (PCD) $\mathcal{P}{\text{full}}$; (b) SOI point cloud after HSV filtering $\mathcal{P}{\text{soi}}$; (c) SOI point cloud after down-sampling and outlier removal; (d) Normal generation; (e) Reconstructed SOI surface using ball pivoting; (f) SOI particles after uniform sampling.
@@ IV Methodology
In this section, we detail our proposed framework for bimanual deformable bag manipulation using a structure-of-interest (SOI) based neural dynamics model. We begin by extracting SOI particles from RGBD data using global particle sampling approach to represent critical bag features, as depicted in Fig. 4(a). A Graph Neural Network (GNN) then models these particles’ dynamics for state prediction (Fig. 4(b)), which informs the Model Predictive Control (MPC) to adjust robotic gripper actions for shaping the SOI (Fig. 4(c)). This integrated approach streamlines the manipulation process, ensuring target SOI configurations are achieved with precision.
Input: SOI observation at time step $t$, $\mathbf{o}^{soi}_{t}$, composed of point cloud data $\mathcal{P}^{1}$, $\mathcal{P}^{2}$, $\mathcal{P}^{3}$ and $\mathcal{P}^{4}$ from four different perspectives.
Output: SOI particle set $\mathbf{P}_{t}$ at time step $t$
1 for each time step $t$ do
2       $\mathcal{P}_{full}$leftarrow point-cloud registration based on $\mathbf{o}^{soi}_{t}$ ;
3       $\mathcal{P}_{soi}$leftarrow extract SOI using $\operatorname{HSV\_filtering}(\mathcal{P}_{full})$;
4       $\mathcal{P}_{soi}$leftarrow downsampling using $\operatorname{VGD}(\mathcal{P}_{soi},v_{s})$ ;
5       $\mathcal{P}_{soi}$leftarrow remove outliers using $\operatorname{SOR}(\mathcal{P}_{soi},k)$ ;
6       $\mathcal{M}_{soi}\leftarrow$ reconstruct surface using $\operatorname{ball\_pivoting}$;
7       $\mathbf{P}_{t}$leftarrow sample particles using $\operatorname{uniform\_sampling}$ $(\mathcal{M}_{soi})$ ;
8       return SOI particle set $\mathbf{P}_{t}$ ;
9
Algorithm 1 Global Particle Sampling